40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Thu Feb 2 15:11:44 2012, Giordon Stark, DailyProgress, Coating Q, NEP of Experimental Set up [FIXED] NoiseMeasurements.pngExperimentalSetup.pngExperimentalSetup.pdf
    Reply  Thu Feb 2 18:59:20 2012, Zach, DailyProgress, Coating Q, front-end noise budget CQ_frontend_electronics_NB_2_2_2102.png
Message ID: 398     Entry time: Thu Feb 2 15:11:44 2012     Reply to this: 399
Author: Giordon Stark 
Type: DailyProgress 
Category: Coating Q 
Subject: NEP of Experimental Set up [FIXED] 

[Giordon, Zach]

 The last post we discussed the noise measurements with a noticeable peak around the 50kHz. This time, we've set the preamp gain to 500 and retook the measurements. The attached image shows the result of said measurements.

One thing we noticed is that the photodetector noise is roughly two orders of magnitude smaller than our best estimates for this. Zach thinks it's not real, I think it's pretty awesome. Here some explanations of the measurements.

  • A-B output means that the output comes from the preamplifier on an AC coupled setting amplified at 500. This either came with the laser on or off. We fixed the range here and then ran a noise floor measurement where we grounded the input.
  • PDA and PDB mean Photodetector A and Photodetector B as listed in the experimental set up drawing in an eLog post some days ago. PDA has a Noise Floor (NFLR) measurement like before as we noticed that the range of it was significantly different from the A-B output range [this is not true of PDB].
  • Theoretical PD and Theoretical Preamp are values highlighted from the rough noise calculations made about a week ago. (Same place as the inclusion of the experimental set up).

I've also noticed that the plot doesn't make complete sense to me. For example, the noise floor of PD1 seems to be higher than PD1 itself - but maybe Zach can re-explain to me the difference between fixed range and auto range here.

For lazy people, I've reattached the experimental set up and noise calculations.

 

Attachment 1: NoiseMeasurements.png  22 kB  | Hide | Hide all
NoiseMeasurements.png
Attachment 2: ExperimentalSetup.png  49 kB  | Hide | Hide all
ExperimentalSetup.png
Attachment 3: ExperimentalSetup.pdf  250 kB  | Hide | Hide all
ExperimentalSetup.pdf ExperimentalSetup.pdf ExperimentalSetup.pdf ExperimentalSetup.pdf ExperimentalSetup.pdf
ELOG V3.1.3-